Grade 12th passModern Physics

A particle oscillates S.H.M with time period 1.2 second.Then Time taken by it in traveling from the extreme to half of amplitude is

Kriti
7 Years agoGrade 12th pass

1 Answer

Eshan
7 Years ago
Dear student,

The equation of SHM can be taken asx=Acos\omega tfor particle starting from extreme position.

\dfrac{A}{2}=Acos\omega t\implies \omega t=\dfrac{\pi}{3}
\implies t=\dfrac{\pi }{3\omega}=\dfrac{T}{6}=0.2s